# Dynamic Target Attack

### v2

@@ 3.3 Optimization — Efficiently Steering Toward the Dynamic Target
so that optimization jointly improves fluency while moving away from refusal-inducing regions in logit space. The overall objective is
$\mathcal{L}_{\text{DTA}}(P,S;r^{*})=\mathcal{L}_{\text{resp}}(P,S;r^{*}_{L})+\lambda\cdot(\mathcal{L}_{\text{flu}}(S)-\mathcal{L}_{\text{rej}}(S)),$ (11)
and we update $S$ by gradient descent (Adam [12] in practice):
$S\leftarrow S-\eta\nabla{S}\mathcal{L}_{\text{DTA}}(P,S;r^{*}).$ (12)
3.4 Iterative Re-Sampling
We note that the first sampled target in a cycle can still be safe. In that case, continued optimization toward this safe target may stall progress and fail to jailbreak the target model. For example, in Figure 3, the first target $T^{1}_{S}$ is more harmful than the initial refusal response $R_{\text{initial}}$ but remains safe, and the first update still yields a safe output $R^{1}_{S}$. After each short optimization round, DTA re-samples from the model’s current conditional distribution given the harmful prompt and the updated suffix. It then selects the most harmful candidate as the next temporary target. This refresh keeps the target anchored to what the model now considers plausible, and it is repeated every cycle rather than focusing to a single target.
Figure 3: Illustration of the core difference between DTA and existing methods.
The brief update toward a model-native target shifts probability distribution away from refusal and reduces the target–distribution discrepancy, which increases the chance that the next re-sampling step surfaces a genuinely harmful, higher-density target. In Figure 3, the second re-sampled target $T^{2}_{S}$ is more harmful than $T^{1}_{S}$. In the second round, DTA optimizes the adversarial suffix $S$ towards $T^{2}_{S}$, successfully inducing the harmful response $R^{2}_{S}$. By alternating short updates with re-sampling, DTA stays aligned with the evolving output space, shortens the optimization path, and converges faster to an effective adversarial suffix. Empirically, this pattern matches Table 4, where increasing the number of re-sampling cycles improves ASR under a fixed compute budget.
@@ 4 Experiments
4.1 Experimental Setup
Target Models. Following prior work [7, 13, 8], we evaluate DTA against five popular open-source LLM families: Llama-3-8B-Instruct (Llama-3) [2], Vicuna-7B-v1.5 (Vicuna) [14], Qwen2.5-7B-Instruct (Qwen2.5) [3], Mistral-7B (Mistral) [15], and Gemma-7B (Gemma) [16]. These models represent diverse architectures and alignment strategies, enabling a comprehensive evaluation and comparison with existing methods. In our default white-box setting, the reference model ($f_{ref}$) is a copy of the target model ($f$). In our black-box setting, we employ two small models, i.e., Llama-3.2-1B-Instruct and Llama-3-8B-Instruct as the surrogate models, to jailbreak two large-scale LLMs, i.e., Llama-3-70B-Instruct and Kimi-K2-Instruct.
Datasets. We conduct experiments on two standard jailbreak benchmarks: AdvBench [6] and HarmBench [17]. To ensure consistent and efficient evaluation while maintaining diversity, we follow previous work [7, 6] and randomly sample a testing subset of 100 prompts from each dataset.
Baselines. Under the white-box setting, we compare DTA against six state-of-the-art white-box methods:COLD-Attack [7], GCG [6], llm-adaptive [10], I-GCG [13], AdvPrefix [8], and RLbreaker [9]. To ensure a fair and rigorous evaluation, all baselines are run with the hyperparameters recommended in their public implementations. Under the black-box setting, we compare DTA against three state-of-the-art black-box methods: ReNeLLM [18], PAP [19], and TAP [20]—again following their default configurations. In addition, we also select two representative white-box methods, COLD-Attack and AdvPrefix, to compare DTA’s performance with their transferability results.
Evaluation metrics. Following previous work [7, 13, 21], we employ a multi-faceted evaluation strategy using three distinct metrics to comprehensively measure attack performance: ${ASR}_{G}$ (judged by GPTFuzzer [22]), ${ASR}_{O}$ (judged by GPT-4), and ${ASR}_{L}$ (judged by Llama-Guard-3-8B [2]). Given a harmful prompt $P$ and the induced response $R$, a judge $\mathcal{J}$ returns an unsafe score. For GPTFuzzer and Llama-Guard-3-8B the score lies in $[0,1]$, denoted as $\mathcal{J}(P,r)$, and we compute $\text{ASR}_{G}$ and $\text{ASR}_{L}$ as
where $\mathbb{I}$ is a indicator function and we set the threshold as $0.5$.
We provide further details in the Appendix.
Implementation Details. In each sampling phase, we sample $N=30$ candidate responses from the target LLM. In the optimization phase, we set the reference model under a standard decoding strategy, e.g., a low decoding temperature of $\tau_{eval}=0.7$ to ensure stable generation. For each of the $M=20$ sampling rounds, the adversarial suffix is optimized for $T=10$ iterations using the Adam optimizer. We apply an early stopping condition, terminating the process for a given prompt if a generated response achieves a harmfulness score above 0.9. All experiments are conducted on two NVIDIA RTX A6000 GPUs, each with 48 GB of memory.
4.2 Main Results under the White-box Setting
Table 1: Comparison of attack success rates (ASRs) achieved by DTA and baselines on five open-sourced LLMs across two datasets. We set M = 20, T = 10 for DTA. We follow the recommended settings for six baselines to achieve their best performance. The red values indicate the highest ASRs.
@@ 4.2 Main Results under the White-box Setting
AdvBench
COLD-Attack 43% 64% 56% 33% 51% 41% 23% 26% 22% 73% 81% 69% 49% 58% 55%
@@
HarmBench
COLD-Attack 28% 42% 46% 18% 29% 37% 23% 37% 31% 42% 58% 50% 38% 46% 44%
@@
Figure 4: Comparison result of DTA and baselines on HarmBench. Dark (Light) bar denotes the average (maximum) ASRs across five target LLMs.
We conduct extensive white-box attack experiments on two widely used benchmarks, AdvBench and HarmBench, across five recent safety-aligned LLMs. In Table 1, we report the detailed ASR results under three evaluation metrics (i.e., GPTFuzzer, GPT-4, and Llama-Guard-3-8B). Figures 5 and 4 show the average (dark bars) and maximum (light bars) ASRs of DTA and existing methods to highlight both effectiveness and stability. Overall, the results clearly show that DTA consistently outperforms existing baselines across two datasets in most cases.
@@
Figure 5: Comparison result of DTA and baselines on AdvBench. Dark (Light) bar denotes the average (maximum) ASRs across five target LLMs.
@@
Beyond the outperforming performance of ASRs, DDTA also exhibits stable behavior across the five target LLM families. On HarmBench, for instance, I-GCG could achieve above 70% $\text{ASR}_{O}$s on Llama-3, Vicuna and Mistral, on Llama-3, Vicuna, and Mistral, yet drops to just 19% on Gemma. A plausible reason is that I-GCG optimizes toward a long, predefined response that typically resides in an extremely low-density region of the model’s conditional output space, which amplifies the target–distribution discrepancy and makes outcomes highly model-dependent. In contrast, under the white-box setting, DTA delivers consistently high performance across diverse LLMs by anchoring its targets in relatively high-density regions via dynamic sampling and iterative re-sampling.
4.3 Results under the Black-box Setting
Under the black-box setting, i.e., only response access to the target model is available, DTA also exhibits its effectiveness in jailbreaking two large LLMs (Llama-3-70B-Instruct and Kimi-K2-Instruct). As shown in Table 2, using Llama-3-8B-Instruct as a surrogate model to jailbreak Llama-3-70B-Instruct, DTA achieves an ASR of 85%, exceeding COLD-Attack (26%) and AdvPrefix (59%) by over 25%. Three black-box jailbreaking methods, ReNeLLM, PAP, and TAP achieve an ASR of 7%, 11%, and 11%, respectively. Against Kimi-K2-Instruct under the same local configuration, DTA reaches 70% ASR, which is competitive with AdvPrefix (72%) and higher than COLD-Attack and three black-box methods. Even with a much smaller local model (Llama-3.2-1B-Instruct), DTA achieves an ASR 30% against Llama-3-70B-Instruct and an ASR of 50% against Kimi-K2-Instruct, respectively. Overall, these results show that DTA remains highly effective in a black-box setting compared to both white-box and black-box methods.
Table 2: Transferability results of DTA across different local LLMs and target (reference) LLMs. In this table, the experiments are conducted on AdvBench and the Attack Success Rates (ASRs) are evaluated by GPTFuzzer (i.e., $\text{ASR}_{G}$). For white-box methods, we employ a local LLM as a surrogate for optimization.
Dataset        Local LLM $\rightarrow$ Target LLM White-box Black-box DTA
COLD-Attack AdvPrefix ReNeLLM PAP TAP
AdvBench
Llama-3.2-1B-it $\rightarrow$ Llama-3-70B-it   7% 53%   7% 11% 11% 30%
Llama-3-8B-it $\rightarrow$ Llama-3-70B-it 26% 59% 85%
Llama-3.2-1B-it $\rightarrow$ Kimi-K2-it   4% 72% 50% 49% 36% 50%
Llama-3-8B-it $\rightarrow$ Kimi-K2-it 20% 72% 70%
4.4 Response diversity
@@ 4.5 Efficiency and Iteration Cost
In Table 3, we present a comparison of the iteration and time costs for DTA and several baselines. The results reveal that existing methods, whether based on fixed-target optimization or more advanced adaptive strategies, often require a substantial number of iterations, typically ranging from 1,000 for methods like GCG and AdvPrefix, to as many as 10,000 for certain adaptive attacks like llm-adaptive-attacks. This high cost stems from the inherent difficulty of forcing a model towards an external, and often unnatural, target distribution.
@@
4.6 Hyper-parameter selection
We conduct an experimental study to explore the impact of the exploration round $M$ and the per-round optimization steps $T$ on DTA’s performance. In Table 4, we report the experimental results of for jailbreaking Llama-3-8B-Instruct on AdvBench while varying the values of $M$ and $T$ under a roughly fixed per-prompt budget $M\times T=200$). The experimental results show that allocating more computational budget to exploration consistently and dramatically improves the ASR.
M = 1 M = 2 M = 4 M = 5 M = 10 M = 20 M = 20 M = 40 M = 100 M = 200
T = 200 T = 100 T = 50 M = 40 T = 20 T = 5 T = 10 T = 5 T = 2 T = 1
34% 41% 65% 70% 85% 89% 92% 95% 97% 100%

### v3

@@ 3.3 Optimization Procedure
so that optimization jointly promotes fluency while pushing the suffix away from refusal-inducing regions in logit space. The overall objective optimized by DTA in each exploration cycle is
$\mathcal{L}_{\mathrm{DTA}}(P,S;r^{*}_{L})=\mathcal{L}_{\mathrm{resp}}(P,S;r^{*}_{L})+\lambda\,\mathcal{L}_{\mathrm{suffix}}(S),$ (12)
where $\lambda>0$ is a hyper-parameter that balances response matching and regularization.
Gradient optimization and projection to discrete tokens.
Given $P$ and $r^{*}_{L}$, we optimize $\mathbf{Z}$ using gradient optimization (following the procedure of COLD-Attack (Guo et al., 2024)). In practice, we use Adam with learning rate $\eta$ and perform $T$ update steps per exploration cycle:
$\mathbf{Z}\leftarrow\mathbf{Z}-\eta\,\nabla_{\mathbf{Z}}\mathcal{L}_{\mathrm{DTA}}(P,S;r^{*}_{L}).$ (13)
Gradients are computed by back-propagating through the softmax-relaxed suffix embeddings and the local LLM $f_{\phi}$. After several steps, we obtain an updated continuous representation of the suffix. To interact with the target LLM and the judge model, DTA periodically projects the continuous suffix back to a discrete sequence, yielding a discrete suffix $S=(s_{1},\ldots,s_{L_{S}})$. This discrete suffix is then used to query the target LLM and will be used to generate candidate responses during the next exploration phase.
@@ 4 Experiments
Table 1: Comparison of attack success rates (ASRs) achieved by DTA and baselines on five recent safety-aligned LLMs across two datasets. We set M = 20, T = 10 for DTA. We adopt the recommended settings for six baselines following existing works (Zhu et al., 2024; Paulus et al., 2024; Chen et al., 2024), and our results align with the findings reported in these studies. We further discuss the results of baselines in Appendix F.1. The red values indicate the highest ASRs.
@@
AdvBench COLD-Attack 43% 64% 56% 33% 51% 41% 23% 26% 22% 73% 81% 69% 49% 58% 55%
@@
HarmBench COLD-Attack 28% 42% 46% 18% 29% 37% 23% 37% 31% 42% 58% 50% 38% 46% 44%
@@
4.1 Experimental Setup
Target models. Following prior white-box work (Guo et al., 2024; Jia et al., 2024; Zhu et al., 2024), we evaluate DTA against five popular open-source LLMs: Llama-3-8B-Instruct (Llama-3) (Dubey et al., 2024), Vicuna-7B-v1.5 (Vicuna) (Chiang et al., 2023), Qwen2.5-7B-Instruct (Qwen2.5) (Yang et al., 2024), Mistral-7B (Mistral) (Jiang et al., 2023), and Gemma-7B (Gemma) (Team et al., 2024). These models represent diverse architectures and alignment strategies, enabling a comprehensive comparison with existing methods. In the black-box setting, we select Llama-3-70B-Instruct and Kimi-K2-Instruct as the target LLMs.
Datasets. We conduct experiments on two standard jailbreak benchmarks: AdvBench (Zou et al., 2023) and HarmBench (Mazeika et al., 2024). To ensure consistent and efficient evaluation while maintaining diversity, we follow previous work (Guo et al., 2024) and randomly sample a testing subset of 100 prompts from each dataset.
Baselines. Under the white-box setting, we compare DTA against six state-of-the-art white-box methods:COLD-Attack (Guo et al., 2024), GCG (Zou et al., 2023), llm-adaptive (Andriushchenko et al., 2025), I-GCG (Jia et al., 2024), AdvPrefix (Zhu et al., 2024), and RLbreaker (Chen et al., 2024). Following previous work (Zhu et al., 2024; Paulus et al., 2024; Andriushchenko et al., 2025), to ensure a fair and rigorous comparison, all white-box baselines are run with the hyper-parameters recommended in their public implementations. Under the black-box setting, we compare DTA against four state-of-the-art black-box methods: GASP (Basani and Zhang, 2024), AdvPrompter (Paulus et al., 2024), PAP (Zeng et al., 2024), and TAP (Mehrotra et al., 2024)—following their default configurations. In addition, we also select two representative white-box methods, COLD-Attack and AdvPrefix, to compare DTA’s performance with their transferability results. We provide baselines’ detailed experimental settings in Appendix G.
Defenses. To evaluate the performance of DTA under jailbreak defenses, we consider two representative mechanisms: Paraphraser (Jain et al., 2023) and SmoothLLM (Robey et al., 2023). For Paraphraser, we use GPT-4 with an instruction-style template to rewrite each adversarial prompt into a paraphrased variant. For SmoothLLM, we perturb adversarial prompts by randomly inserting additional characters as described in the original paper. More details are provided in Appendix C.4.
Evaluation metrics. Following previous work (Guo et al., 2024; Jia et al., 2024; Paulus et al., 2024), we employ a multi-faceted evaluation strategy using three distinct metrics to comprehensively measure attack performance: ${ASR}_{G}$ (judged by GPTFuzzer (Yu et al., 2024)), ${ASR}_{O}$ (judged by GPT-4), and ${ASR}_{L}$ (judged by Llama-Guard-3-8B (Dubey et al., 2024)). Given a harmful prompt $P$ and the induced response $R$, a judge $\mathcal{J}$ returns an unsafe score. For GPTFuzzer and Llama-Guard-3-8B the score lies in $[0,1]$, denoted as $\mathcal{J}(P,r)$, and we compute $\text{ASR}_{G}$ and $\text{ASR}_{L}$ as
where $\mathbb{I}$ is a indicator function and we set the threshold as $0.5$ ($\geq 0.5$ denotes harmfulness).
We provide further details in Appendix C.
Implementation details. During exploration, we sample $N=30$ candidates from the target LLM. In the optimization phase, we set the local LLM under a standard decoding strategy, e.g., a low decoding temperature of $\tau_{eval}=0.7$ to ensure stable generation. For each of the $M=20$ exploration cycles, the adversarial suffix is optimized for $T=10$ iterations using the Adam optimizer. We apply an early stopping condition, terminating the process for a given prompt if a generated response achieves a harmfulness score above 0.9. All experiments are conducted on two NVIDIA RTX A6000 GPUs, each with 48 GB of memory.
4.2 Main Results under the White-box Setting
Figure 4: Comparison results of DTA and baselines on HarmBench (left) and AdvBench (right). Dark (Light) bars denote the average (maximum) ASRs across five target LLMs.
We conduct extensive white-box attack experiments on two widely used benchmarks, AdvBench and HarmBench, across five recent safety-aligned LLMs. In Table 1, we report the detailed ASR results under three evaluation metrics (i.e., GPTFuzzer, GPT-4, and Llama-Guard-3-8B). Figure 4 shows the average (dark bars) and maximum (light bars) ASRs of DTA and existing methods to highlight both effectiveness and stability. Overall, the results clearly show that DTA consistently outperforms existing baselines across two datasets in most cases.
@@ 4.2 Main Results under the White-box Setting
Beyond the outperforming performance of ASRs, DTA also exhibits stable behavior across the five target LLM families. On HarmBench, for instance, I-GCG could achieve above 70% $\text{ASR}_{O}$s on Llama-3, Vicuna and Mistral, on Llama-3, Vicuna, and Mistral, yet drops to just 19% on Gemma. A plausible reason is that I-GCG optimizes toward a long, predefined response that typically resides in an extremely low-density region of the model’s conditional output space, which amplifies the target–distribution discrepancy and makes outcomes highly model-dependent. In contrast, under the white-box setting, DTA delivers consistently high performance across diverse LLMs by anchoring its targets in relatively high-density regions via dynamic sampling and iterative re-sampling. We further discuss the results of baselines in Appendix F.1.
Table 2: Results under the black-box setting. We employ Llama-3-8B-it as a surrogate to jailbreak the target LLM (Llama-3-70B-it/Kimi-K2-it). For example, AdvPrompter and GASP use Llama-3-8B-it as a surrogate to craft adversarial prompts, then we use these prompts to query the target LLMs. The results indicate these methods’ low performance on transfer attacks.
Target LLM COLD AdvPrefix AdvPrompter GASP TAP DTA
Llama-3-70B-it 26% 59% 35% 12% 11% 85%
Kimi-K2-it 20% 72%   5% 23% 36% 70%
Average 23.0% 65.5% 20.0% 17.5% 23.5% 77.5%
4.3 Main Results under the Black-box Setting
Table 2 reports the black-box transferability results of DTA, where a local surrogate model (Llama-3-8B-it) is used to craft adversarial suffixes that are then directly transferred to stronger target LLMs (Llama-3-70B-it and Kimi-K2-it) without any access to their parameters or gradients. As shown in the table, DTA significantly outperforms prior baselines across both target models. Specifically, DTA achieves an average ASR of 77.5%, substantially higher than existing baselines, e.g., COLD-Attack (23.0%), AdvPrefix (65.5%).
These results indicate that the adversarial suffixes optimized by DTA generalize effectively across model scales and architectures, despite being trained solely on a smaller surrogate model. In contrast, existing baselines exhibit limited transferability, often suffering from severe performance degradation in the black-box setting. This suggests that DTA’s dynamic-target optimization objective captures more model-agnostic harmful response patterns, rather than overfitting to surrogate-specific decoding behaviors.
4.4 Ablation Study
We conduct ablation studies on three key hyperparameters of DTA: the forward truncation length $L$, the exploration temperature $\tau_{\text{search}}$, and the sampling count per exploration cycle $N$. More results are provided in Appendix F.
Figure 5: Ablation study on the sampling count.
Sampling count $N$.
As shown in Figure 5, we vary $N$ from $1$ to $50$ on AdvBench against Llama-3-8B-Instruct and observe that the final $\text{ASR}_{G}$ remains in a narrow band of roughly $87\%\text{--}94\%$ across all settings. Even with N=1N{=}1, DTA already achieves $90\%$ $\text{ASR}_{G}$, while a moderate choice N=30N{=}30 reaches the highest $\text{ASR}_{G}$ and yields the fastest early successes (over $80\%$ of successful jailbreaks within the first 5 exploration cycles). Larger $N$ increases per-cycle cost without bringing consistent ASR gains, indicating that DTA is not overly sensitive to the sampling count and that small-to-moderate $N$ already provides sufficient exploration.
Figure 6: Ablation study on truncation length.
Truncation length $L$.
Figure 5 shows the ablation results on the truncation length $L$ and the results indicate that DTA remains highly effective across a wide range of $L$: even L=5L{=}5 attains $87\%$ $\text{ASR}_{G}$, and increasing $L$ to 10102020 already achieves $93\%\text{--}94\%$. For moderate lengths $L\in[20,100]$, the ASR fluctuates mildly around $90\%$, suggesting that DTA does not rely on matching very long targets. Very large $L$ values bring no consistent benefit and can slightly degrade performance while increasing computational cost. Based on this trade-off, we fix L=20L{=}20 in the main experiments.
Figure 7: Ablation study on truncation length.
Exploration temperature $\tau_{\text{search}}$.
We further sweep $\tau_{\text{search}}$ from $0.1$ to $4.0$ and the results are shown in Figure 7. The $\text{ASR}_{G}$ stays high (around $88\%\text{--}94\%$) throughout this range: even an extremely low temperature $\tau_{\text{search}}{=}0.1$ yields $88\%$ $\text{ASR}_{G}$, showing that DTA can effectively increase the probability of harmful completions even under conservative sampling. Around $\tau_{\text{search}}\!\approx\!2.0$, the ASR stabilizes in the $92\%\text{--}94\%$ range, and further increasing the temperature does not lead to systematic improvements. In our main experiments, we therefore set $\tau_{\text{search}}{=}2.0$ as a balanced choice that provides sufficient diversity while maintaining stable performance. Table 14 shows some examples of DTA sampled and generated responses under different exploration temperatures.
4.5 Results under defense mechanisms
Table 3: Results under defense methods. Overall $\text{ASR}_{O}$ of COLD-Attack and DTA on AdvBench against Llama-3-8B-it.
Method COLD-Attack DTA
No defense 64% 98%
Paraphrase 25% 46%
SmoothLLM 56% 89%
We further evaluate the robustness of DTA against two recently proposed jailbreak defenses, paraphrase-based input obfuscation and SmoothLLM, on AdvBench with Llama-3-8B-it (Table 3). Without any defense, DTA already achieves a near-perfect overall success rate of 98%, substantially higher than COLD-Attack (64%). Under the paraphrase defense, both methods suffer a drop in performance, but DTA still maintains 46% ASRO compared to 25% for COLD-Attack. When SmoothLLM is applied, DTA remains highly effective with 89% ASRO, whereas COLD-Attack reaches only 56%. These results indicate that, across both types of defenses, DTA consistently outperforms COLD-Attack and exhibits markedly stronger robustness under defense, while retaining very high attack success in the undefended setting.
@@ 4.6 Efficiency and Iteration Cost
In Table 4, we present a comparison of the iteration and time costs for DTA and several baselines. The results reveal that existing methods often require a substantial number of iterations, typically ranging from 1,000 for methods like GCG, to as many as 10,000 for certain adaptive attacks like llm-adaptive-attacks. This high cost stems from the inherent difficulty of forcing a model towards an external, and often unnatural, target distribution.